# Identifying treatment response subgroups in observational time-to-event data

### v3

@@ Abstract
Identifying patient subgroups with different treatment responses is an important task to inform medical recommendations, guidelines, and the design of future clinical trials. Existing approaches for subgroup analysis primarily rely on Randomised Controlled Trials (RCTs), in which treatment assignment is randomised. RCTs’ patient cohorts are often constrained by cost, rendering them not representative of the heterogeneity of patients likely to receive treatment in real-world clinical practice. When applied to observational studies, subgroup analysis approaches suffer from significant statistical biases particularly because of the non-randomisation of treatment. Our work introduces a novel, outcome-guided method for identifying treatment response subgroups in observational studies. Our approach assigns each patient to a subgroup associated with two time-to-event distributions: one under treatment and one under control regime. It hence positions itself in between individualised and average treatment effect estimation. The assumptions of our model result in a simple correction of the statistical bias from treatment non-randomisation through inverse propensity weighting. In experiments, our approach significantly outperforms the current state-of-the-art method for outcome-guided subgroup analysis in both randomised and observational treatment regimes.
@@ 1 Introduction
Understanding heterogeneous therapeutic responses between patient subgroups is the core of treatment guidelines and the development of new drugs. Identifying such subgroups is valuable to inform clinical trials by identifying subgroups not responding to existing drugs, and to direct healthcare resources to those who might benefit most, and away from those who may be most harmed (Foster et al., 2011). For instance, subgroup analysis in the BARI trial of patients with coronary artery disease supported the use of coronary artery bypass graft over percutaneous interventions for patients with diabetes, and the opposite for patients without diabetes (Investigators, 1996), shaping subsequent guidelines. Figure 1 illustrates this idea: two groups may present opposite treatment responses, and would benefit from different recommendations. Our work aims to uncover such subgroups.
Randomised controlled trials (RCTs) remain the gold standard for identifying subgroups of treatment effects. In RCTs, patients are randomly assigned to control or treated groups, allowing researchers to assess the impact of an intervention or treatment. However, RCTs are costly and time-consuming, with restricted patient cohorts unrepresentative of the real-world diversity of patients who may receive treatment and treatment strategies (Hernán & Robins, 2016).
Our work introduces a novel approach that diverges from traditional RCT-based methodologies (Nagpal et al., 2022a; 2023) by leveraging routinely collected observational data to identify patient subgroups. Observational studies encompass larger and more diverse cohorts reflective of real-world practices, offering the potential to uncover subgroups of treatment responses, that could be missed in RCTs. Prior works (Bica et al., 2020; Curth et al., 2021; Louizos et al., 2017) have leveraged observational data while addressing biases from non-random treatment assignments (Benson & Hartz, 2000; Hernán & Robins, 2010; Hernán, 2018). However, this body of literature primarily focuses on estimating (i) averaged treatment effects at the population level or (ii) individualised treatment effects, thereby overlooking the identification of treatment effect subgroups.
Addressing this critical gap in the literature, our work uncovers patient subgroups with distinct treatment responses using observational data. Unlike previous methods relying on RCTs, we introduce a mixture of neural networks as an extension to traditional outcome-guided models to uncover subgroups delineated by non-linear, higher-dimensional combinations of available covariates, without requiring parametrisation of the time-to-event distributions or treatment responses.
@@
Figure 1: Subgroup treatment effect discovery in time-to-event observational data. Our method identifies subgroups of patients with similar treatment responses to guide clinical practice and design clinical trials. Our method simultaneously models the treatment effect and identifies subgroup while addressing censoring and treatment non-randomisation.
Contributions.
Our work introduces the first neural network approach to simultaneously estimate treatment effects and discover subgroups in observational settings, addressing both censoring and non-random treatment assignments without parameterising the survival distribution or treatment effects. Section 2 formalises the problem of treatment subgroup discovery and non-randomisation correction. Then, we introduce the proposed monotonic neural network implementation in Section 2. Finally, we evaluate the proposed methodology on two synthetic datasets in Section 3, an extensive set of alternative settings in App. B and a real-world example described in App. C and compare it with related approaches reviewed in Section 4.
@@ 2.1 Problem setup
Our goal is to uncover subgroups with similar treatment responses, guided by the observed times of occurrence of an outcome of interest. Patients are assigned to subgroups based on their covariates at treatment time. Each subgroup is characterised by two distributions, known as survival functions: one under treatment and one under control regimes. These distributions characterise the probability of observing the event of interest after a given time under a given treatment.
Consider the random variables associated with observed covariates $X$, an indicator identifying whether the event of interest, in our analysis death, was observed $D$, and the observed time of event $T$. Formally, we define the latter random variables as $T:=\min(C,T^{\prime})$ and $D:=\mathbbm{1}(C>T^{\prime})$, where $C$ is random variable of the (right)-censoring time, i.e. the time at which the patient left the study before experiencing the event of interest, and $T^{\prime}$ is the partially observed random variable associated with the time of the event of interest if there was no censoring. When a patient is censored, $T=C$ and $D=0$, otherwise the event is observed and $T=T^{\prime}$ and $D=1$. To estimate the associated survival likelihood, we further assume, as commonly done in survival analysis, that the censoring time $C$ is not informative for the time of the event of interest $T^{\prime}$.
Assumption 1 (Non-informative censoring).
Central to our problem is the additional variable $Z$ associated with the latent, unobserved subgroup membership. Following Jeanselme et al. (2022), we aim to identify a pre-specified $K\in\mathbb{N}$ number of subgroups [^1] As a pre-specified number of subgroups may be a limitation in a real-world setting where we do not know the underlying grouping structure, we explore how to select this parameter based on the likelihood of the predicted outcomes in Section 3.. We assume that $Z$ depends on $X$ and influences $T^{\prime}$. As we are interested in recovering the survival functions associated with the latent subgroups, we ignore the potential dependence of $T^{\prime}$ on $X$, considering the individual survival distribution as a mixture of the different subgroups.
Assumption 2 (Mixture modelling).
The event time $T^{\prime}$ is completely determined by the patient’s group membership $Z$. Formally, $T^{\prime}\perp\!\!\!\perp X\mid Z$.
While this assumption may hurt individual performance as individual covariates do not directly inform individual survival but group membership only, it improves the interpretability of the model by separating subgroup membership from survival profiles. In conclusion, we aim to recover $Z$ and the survival distributions associated with each subgroup from the observed $X$, $T$ and $D$.
Given the previously described dependencies, the expected survival $S$ for a given patient with covariates $x$ is the survival when marginalising over the different subgroups:
$S(t\mid X=x):=\mathbb{P}(T^{\prime}\geq t\mid X=x)=\sum_{k=1}^{K}\mathbb{P}(Z=k\mid X=x)\mathbb{P}(T^{\prime}\geq t\mid Z=k)$ (1)
Note that the last term is linked to the quantity commonly used in the survival literature: $\Lambda_{k}(t):=-\log\mathbb{P}(T^{\prime}\geq t\mid Z=k)$, the group-specific cumulative hazard, with derivative with respect to time $\lambda_{k}(t)$, the group-specific instantaneous hazard, corresponding to the increase in the probability of observing the event of interest given survival until that time $t$. Estimating the two probability distributions $\mathbb{P}(Z=k\mid X=x)$ and $\mathbb{P}(T^{\prime}\geq t\mid Z=k)$ is our core interest.
Likelihood-based model fitting.
One can maximise the log-likelihood associated with the observed realisations $\{x_{i},t_{i},d_{i}\}$ of the random variables $X,T,D$ for patient $i\in[\![1,N]\!]$ with $N$ the total number of patients (we omit $i$ where redundant) to estimate the survival function (Eq. equation 1). Under Assumption 1, Equation equation 2 describes this factual [^2] Referring to the likelihood of the observed (rather than counterfactual) realisations. log-likelihood.
$l_{F}(\theta)$ $=\sum_{i,d_{i}=1}\log\left(-\frac{\partial S_{\theta}(u\mid x_{i})}{\partial u}\bigg{|}_{u=t_{i}}\right)+\sum_{i,d_{i}=0}\log S_{\theta}(t_{i}\mid x_{i})$ (2)
where $\theta$ is the set of parameters characterising the estimated survival function $S_{\theta}$. Here, patients with an observed event ($d_{i}=1$) contribute the negative derivative of the survival function to the likelihood (see Jeanselme et al. (2022) for derivation). Each censored patient ($d_{i}=0$) contributes the probability of not observing an event before $t_{i}$, i.e., $S_{\theta}(t_{i}\mid x_{i})$, under assumption 1 of independence between the censoring time and the event of interest conditional on the covariates.
Latent treatment effect subgroups.
@@
Consider now the binary treatment variable $A$. Patients either receive the treatment ($A=1$) or they do not ($A=0$). Therefore, following the potential outcomes formulation, we consider the time of events under treatment $T^{\prime}_{1}$ and under the control regime $T^{\prime}_{0}$. The central challenge is that one can not observe both $T^{\prime}_{0}$ and $T^{\prime}_{1}$, but only $T^{\prime}$, the random variable associated with the event time under the observed treatment regime in the absence of censoring, which is equal to $A\cdot T^{\prime}_{1}+(1-A)\cdot T^{\prime}_{0}$ under assumption 4 described below. Figure 2 describes the dependencies between the previously described random variables.
A critical assumption of our proposed setting is that the subgroups we aim to identify do not influence treatment if we know a patient’s covariates, formalised as follows:
Assumption 3 (Unknown latent groups).
The treatment assignment $A$ is independent of the subgroup membership $Z$ given the covariates $X$. Formally, $A\perp\!\!\!\perp Z\mid X$.
This assumption relaxes the traditional assumption of treatment randomisation. Note that this assumption remains realistic, as (i) clinicians base treatment recommendations on the patients’ covariates, and (ii) if subgroups were known, one would not need the proposed methodology to uncover novel subgroups.
Figure 2: Graphical representation between covariates ($X$) and outcomes ($T,D$). Realisations of dashed variables are unobserved, while $X$, $A$, $T$ and $D$ are observed.
Further, the following three assumptions are necessary to estimate treatment effects and are common in the causal literature.
@@
A patient’s observed event time is the potential event time associated with the observed treatment. Formally, this means $T^{\prime}:=A\cdot T^{\prime}_{1}+(1-A)\cdot T^{\prime}_{0}$ where $T^{\prime}$ is the observed event time and $(T^{\prime}_{0},T^{\prime}_{1})$ are the potential event times under the treatment $A$.
@@
Under the Assumptions 4, 5, and 6, existing works often focus on estimating the Individualised Treatment Effect (ITE, here denoted as $\tau$), defined as the difference between survival under the two potential treatment regimes given the covariates $x$ (see App. A.1 for derivation):
$\tau(t,x)=\mathbb{P}(T^{\prime}\geq t\mid A=1,X=x)-\mathbb{P}(T^{\prime}\geq t\mid A=0,X=x)$ (3)
Estimating this quantity requires accurately modelling the survival distributions under the two treatment regimes for all $x$ from observed data. This estimation would be straightforward if one could access the observed survival times for all patients under both treatment regimes. In this case, one would estimate the survival using the event times under $A=1$ and under $A=0$. The key challenge is that the counterfactual survival outcome is not observed: if a patient receives the treatment, we do not observe its outcome under no treatment, and vice versa.
In RCTs, treatment $A$ and covariates $X$ are by design independent. Hence, the subset of the patients receiving either treatment regime is representative of the overall population. Relying on this observed subset to estimate the survival distribution under each regime is a valid estimate for the population. In other words, maximising the factual likelihood to estimate survival under both treatment regimes results in a valid estimate. However, covariates and treatment are not independent in observational studies in which treatment recommendations depend upon the observed covariates. Formally, $\mathbb{P}(A\mid X)\neq\mathbb{P}(A)$ in general. For example, clinicians might recommend more aggressive treatment to patients in more severe conditions. This absence of randomisation results in a covariate shift between the treated and non-treated populations (Curth et al., 2021) as their covariate distributions differ (Bica et al., 2021). In this setting, estimating survival by maximising the factual likelihood is no-longer enough to estimate the counterfactual survival distribution.
Under assumption 6, approaches such as re-weighting (Shimodaira, 2000), or penalisation on the dissimilarity of learnt representations (Johansson et al., 2016; Shi et al., 2019), or a combination of these approaches (Curth et al., 2021; Hassanpour & Greiner, 2019; Shalit et al., 2017) are remedies to estimate the counterfactual likelihood from the observed data by reducing the difference between the two treatment regimes’ populations when modelling the outcome of interest. Specifically, Shalit et al. (2017) demonstrate that the negative log likelihood is upper-bounded by observable quantities. Here we extend their notations to the survival setting as
$-l(\theta)\leq-l_{F}^{*}(\theta)+\gamma\cdot\text{IPM}(q_{\Phi}^{A=0},q_{\Phi}^{A=1}),$ (4)
where $l$ is the log-likelihood consisting of both factual and counterfactual log-likelihoods, $l_{F}^{*}$ is the factual log-likelihood (defined in equation 2) weighted with an inverse propensity of treatment weighting for each patient, $q_{\Phi}^{A=a}=q(\Phi(X)\mid A=a)$ is the density function associated with the transformation $\Phi$ of the covariates $X$, the Integral Probability Metric (IPM) is a distance between distributions, and $\gamma$ is a positive constant.
Shalit et al. propose to use this bound to train a neural network to estimate treatment effects. In this context, $\Phi$ is an inner representation of the network and the IPM regularisation renders this embedding similar between the treated and untreated populations. This embedding is then used to estimate the survival function under both treatment regimes. The difference between the estimated survival functions is then an accurate estimate of the treatment effect as the IPM penalisation corrects for the shift resulting from treatment non-randomisation.
Note that, under the dependencies we assume, only the subgroup assignment ($\mathbb{P}(Z\mid X)$) depends on the covariates $X$ and could, consequently, play the role of the transformation $\Phi$. However, treatment is assumed independent of the group membership under assumption 3, meaning that treatment rate does not differ between subgroups. This assumption results in the regularisation term being null, i.e. $\text{IPM}(q_{Z\mid X}^{A=0},q_{Z\mid X}^{A=1})=0$ (see App. A.2 for derivation). This assumption is consequently necessary to correctly estimate the subgroups’ treatment effect, without penalising the clustering structure.
As a consequence, our work focuses on the first term of the upper bound in equation 4 by re-weighting the factual likelihood with the patient’s propensity score (Hassanpour & Greiner, 2019; Shalit et al., 2017). Using an estimator $\hat{p}_{A}(x):=\mathbb{P}(A=1\mid X=x)$ of the propensity score, we use a truncated propensity weighting (Austin & Stuart, 2015) scheme to avoid unstable weights in which the weights $w_{i}$ are defined as
$w_{i}^{-1}=\begin{cases}0.05&\text{ if }\hat{p}_{A}(x_{i})<0.05\\ 0.95&\text{ if }\hat{p}_{A}(x_{i})>0.95\\ a_{i}\cdot\hat{p}_{A}(x_{i})+(1-a_{i})\cdot(1-\hat{p}_{A}(x_{i}))&\text{ otherwise},\end{cases}$ (5)
where $i$ is the patient index, and $a_{i}$ is the realisation of $A$ for patient $i$. Using the factual likelihood from equation 2 and the weights $w_{i}$ defined in equation 5, we derive the upper-bound of the Negative Log-Likelihood (NLL), later used to train our model, as
$\sum_{i,d_{i}=1}w_{i}\log\left(-\frac{\partial S_{\theta}(u\mid x_{i},a_{i})}{\partial u}\bigg{|}_{u=t_{i}}\right)+\sum_{i,d_{i}=0}w_{i}\log S_{\theta}(t_{i}\mid x_{i},a_{i})$ (6)
@@
Figure 3: Causal Survival Clustering architecture. Latent parameter $l_{k}$ characterising the subgroup $k$ is inputted in the monotonic network $M$ to estimate the cumulative hazard $\Lambda_{k}$ under both treatment regimes. $G$ assigns the probability to belong to each subgroup given the patient’s covariate(s) $x$. To tackle the challenge of treatment assignment bias, the network $W$ estimates the treatment propensity used to weigh the training likelihood. In this context, patient survival is the average of the weighted survival distributions across subgroups under the given treatment.
2.2 Estimating the quantities of interest with neural networks

### v4

@@ Abstract
Identifying patient subgroups with different treatment responses is an important task to inform medical recommendations, guidelines, and the design of future clinical trials. Existing approaches for treatment effect estimation primarily rely on Randomised Controlled Trials (RCTs), which are often limited by insufficient power, multiple comparisons, and unbalanced covariates. In addition, RCTs tend to feature more homogeneous patient groups, making them less relevant for uncovering subgroups in the population encountered in real-world clinical practice. Subgroup analyses established for RCTs suffer from significant statistical biases when applied to observational studies, which benefit from larger and more representative populations. Our work introduces a novel, outcome-guided, subgroup analysis strategy for identifying subgroups of treatment response in both RCTs and observational studies alike. It hence positions itself in-between individualised and average treatment effect estimation to uncover patient subgroups with distinct treatment responses, critical for actionable insights that may influence treatment guidelines. In experiments, our approach significantly outperforms the current state-of-the-art method for subgroup analysis in both randomised and observational treatment regimes.
Machine Learning, ICML
@@ 1 Introduction
Understanding heterogeneous therapeutic responses among patient subgroups is central to developing clinical guidelines and new treatments. Identifying such subgroups is valuable to inform the design of future clinical trials and to direct healthcare resources to those most likely to benefit, and away from those at greatest risk of harm (Foster et al., 2011). A striking example comes from the BARI trial on coronary artery disease, which suggested that patients with diabetes should receive coronary artery bypass grafts rather than percutaneous interventions, whereas patients without diabetes benefited from the opposite strategy (Investigators, 1996), an observation that shaped subsequent guidelines. Figure 1 illustrates this concept: two groups with opposing treatment responses may require distinct therapeutic recommendations. Our work aims to uncover such subgroups in observational, time-to-event data.
Randomised controlled trials (RCTs) remain the gold standard for studying heterogeneous treatment effects. By randomly assigning patients to control or treated arms, RCTs eliminate potential confounders when assessing a treatment’s impact on outcomes. However, RCTs are often costly, time-consuming, and restricted to specific patient cohorts, which are likely not to reflect the full spectrum of patients seen in clinical practice (Hernán & Robins, 2016, 2010; Cole & Hernán, 2008). Trial findings may thus not generalise to real-world settings, where treatment choice depends on clinical judgment and patient preferences, among other variables that may remain unobserved.
@@
Figure 1: Subgroup treatment effect discovery in time-to-event observational data. Our work aims to identify subgroups of patients with similar treatment responses to guide clinical practice and design clinical trials. Our method simultaneously models the treatment effect and identifies subgroup while addressing treatment non-randomisation and censoring.
To address these limitations, researchers increasingly leverage large-scale observational datasets which capture diverse, real-world populations. Although observational data offer the potential to identify subgroups that RCTs might miss, they also pose challenges due to non-random treatment assignments and confounding (Benson & Hartz, 2000; Hernán & Robins, 2010; Hernán, 2018). A robust methodological literature in causal inference has emerged to tackle these challenges. Techniques such as inverse probability weighting (Cole & Hernán, 2008), marginal structural models (Robins et al., 2000), and doubly robust estimators (Bang & Robins, 2005) allow for more accurate causal effect estimation under appropriate assumptions.
Prior works in machine learning (ML) have extended these ideas to estimate treatment effects from observational data (Bica et al., 2020; Curth et al., 2021; Louizos et al., 2017). However, most methods focus on (i) average treatment effects at the population level or (ii) individualised treatment effects. These approaches overlook the utility of identifying treatment-effect subgroups —groups with consistently distinct responses to a given intervention. Detecting such subgroups is crucial in translating treatment effect estimates into actionable insights for clinical decision-making and resource allocation.
We address this gap by introducing a novel framework that uncovers patient subgroups with distinct treatment responses using observational time-to-event data. Extending the rich tradition of outcome-guided modelling (Foster et al., 2011) to a flexible, neural network–based architecture. Crucially, our approach does not require explicit parametrisation of time-to-event distributions or treatment effects; it instead leverages a mixture of monotonic neural networks with correction for non-random treatment assignment. In summary, our contributions are:
  • Neural network-based treatment subgroup identification. Our work introduces the first neural network architecture to simultaneously estimate and identify subgroups of treatment effects in observational settings, accommodating both censoring and non-random treatment assignments without parameterising the treatment effects.
  • Novel formalisation. In section 2, we review existing approaches to identify subgroups of treatment effects. Section 3 formalises the problem of subgroup discovery and our strategy for minimising biases arising from non-random treatment assignment.
  • Extensive application on survival time-to-event data. In section 3, we introduce the proposed monotonic neural network implementation, and evaluate its performance on two synthetic datasets in Section 4, an extensive set of alternative settings in App. B and a real-world example described in App. C, benchmarking against several established approaches.
2 Related work
While survival subgrouping has been proposed through mixtures of distributions (Nagpal et al., 2021a, b; Jeanselme et al., 2022) to identify different phenotypes of patients, the ML literature on phenotyping treatment effects with time-to-event outcomes remains sparse. The current focus is on estimating population or individual treatment effects in the static (Shalit et al., 2017; Johansson et al., 2016; Zhang et al., 2020) and survival settings (Curth et al., 2021). While these approaches help to understand the population or individualised treatment effects, they do not identify subgroups that may benefit or be harmed by treatment. Identifying such groups aligns with, and is thus more useful for, drafting medical guidelines that direct treatment to those subgroups most likely to benefit from it.
Tree-based approaches were proposed to address the limitations of step-wise approaches by jointly discovering subgroups and modelling the treatment effect. Instead of traditional splits on the observed outcomes, these causal trees aim to discover homogeneous splits regarding covariates and treatment effects. Su et al. (2009) introduce a recursive population splitting based on the average difference in treatment effect between splits. Athey & Imbens (2015, 2016) improve the confidence interval estimation through the honest splitting criterion, which dissociates the splitting from the treatment effect estimation. Wager & Athey (2018) agglomerate these causal trees into causal forests for improved ITE estimation. Each obtained split in the decision tree delineates two subgroups of treatment effect (Lipkovich et al., 2011; Loh et al., 2015). Alternatively, McFowland III et al. (2018) propose pattern detection and Wang & Rudin (2022), causal rule set learning to uncover these subgroups. However, all these approaches rely on a local optimisation criterion (Lipkovich & Dmitrienko, 2014) and greedy split exploration. Recently, Nagpal et al. (2020) addressed the local optimisation by constraining the treatment response to a linear form in a mixture of Cox models.
Previous approaches uncover subgroups of treatment effect but consider RCTs with binary outcomes, not the observational setting with survival outcomes that our work explores. At the intersection with survival analysis, Zhang et al. (2017) extend causal trees to survival causal trees, modifying the splitting criterion by measuring the difference in survival estimates between resulting leaves. Similarly, Hu et al. (2021) propose Bayesian additive models and Zhu & Gallego (2020) propose a step-wise approach with propensity weighting to study observational data. Closest to our work, (Jia et al., 2021; Nagpal et al., 2023) propose to uncover subgroups within RCTs with survival outcomes. Jia et al. (2021) propose a mixture of treatment effects characterised by Weibull distributions trained in an expectation-maximisation framework. Similarly, Nagpal et al. (2023) stratify the population into three groups: non-, positive- and negative responders to treatment. An iterative Monte Carlo optimisation is used to uncover these subgroups, characterised by a Cox model with a multiplicative treatment effect. As demonstrated in our work, this step-wise optimisation may be limiting, and the assumption of RCTs renders the model less relevant in observational data.
@@ 3.1 Problem setup
Our goal is to uncover a pre-specified $K\in\mathbb{N}$ number of subgroups [^1] As a pre-specified number of subgroups may be a limitation in a real-world setting where we do not know the underlying grouping structure, we explore how to select this parameter based on the likelihood of the predicted outcomes in Section 4. with distinct treatment responses, guided by the observed times of occurrence of an outcome of interest. To this end, our model assigns patients to latent subgroups based on their covariates at the treatment decision time. For each subgroup, we estimate the probability of not observing the event of interest given time under treatment and under control regimes. These two distributions are known as survival functions, and their difference corresponds to the estimated treatment effect for a given group.
@@
Formally, consider the random variables associated with observed covariates $X$, the assigned binary treatment $A$, and the observed event time $T^{\prime}$. Following the potential outcomes formulation, $T^{\prime}=A\cdot T^{\prime}_{1}+(1-A)\cdot T^{\prime}_{0}$ under consistency (Assumption 1) where $T^{\prime}_{1}$ is the potential event time under treatment and $T^{\prime}_{0}$, under the control regime.
@@
A patient’s observed event time is the potential event time associated with the observed treatment. Formally, this means $T^{\prime}=A\cdot T^{\prime}_{1}+(1-A)\cdot T^{\prime}_{0}$ where $T^{\prime}$ is the observed event time and $(T^{\prime}_{0},T^{\prime}_{1})$ are the potential event times under the control and treatment regimes respectively.
Central to our problem is the latent, unobserved, subgroup membership $Z$. As formalised in Assumption 2, we assume that the variable $Z$ mitigates the dependence between the covariates $X$ and the observed event time $T^{\prime}$.
Assumption 2 (Mixture mitigation).
The event time $T^{\prime}$ is determined by the patient’s group membership $Z$. Formally, $T^{\prime}\perp\!\!\!\perp X\mid Z$.
In this setting, our interest lies in estimating the Subgroup Average Treatment Effect (SATE) for group $k$:
$\tau_{k}(t)$ $=\mathbb{P}(T^{\prime}\geq t\mid A=1,Z=k)$
$-\mathbb{P}(T^{\prime}\geq t\mid A=0,Z=k)$ (1)
This focus differs from the literature interest in estimating Individualised Treatment Effects (ITE), defined as follows (see App. A.1 for derivation from the potential outcomes formalisation):
$\tau(t,x)$ $=\mathbb{P}(T^{\prime}\geq t\mid A=1,X=x)$
$-\mathbb{P}(T^{\prime}\geq t\mid A=0,X=x)$
$=\sum_{k}\tau_{k}(t)\mathbb{P}(Z=k\mid X=x)$ (Under Asm. 2)
Estimating the previous quantities requires the accurate modelling of the survival distributions under the two treatment regimes. The central challenge in this estimation is that the counterfactual survival outcome is unobserved: if a patient receives the treatment, we do not observe its outcome under no treatment, and vice versa.
The absence of treatment randomisation in observational studies, e.g. clinicians might recommend more aggressive treatment to patients in more severe conditions, results in a covariate shift between the treated and non-treated populations (Curth et al., 2021) as their covariate distributions differ (Bica et al., 2021). This prohibits the estimation of the survival functions through the maximisation of the observed outcomes likelihood.
Under Assumptions 3 and 4, remedies such as re-weighting (Shimodaira, 2000), or penalisation on the dissimilarity between treatment regimes (Johansson et al., 2016; Shi et al., 2019), or a combination of these approaches (Curth et al., 2021; Hassanpour & Greiner, 2019; Hernán & Robins, 2006; Shalit et al., 2017) allows to estimate both factual and counterfactual likelihood from the observed data by penalising the difference between the treated and non-treated populations. In our work, we adopt an Inverse Propensity Weighting (IPW) of the factual likelihood to estimate the overall likelihood, consisting in weighting each observation by the inverse probability of receiving treatment.
@@
To express the factual likelihood, one needs to tackle a final challenge in time-to-event data: censoring. Patients may not observe the outcome of interest during the study period or exit the study for reasons independent of the outcome of interest. Formally, instead of observing $T^{\prime}$, one observes an indicator identifying whether the event of interest was observed $D$, and the associated observed time of event $T$, with $T:=\min(C,T^{\prime})$ and $D:=\mathbbm{1}(C>T^{\prime})$, where $C$ is the random variable of the (right)-censoring time. When a patient is censored, $T=C$ and $D=0$; otherwise, the event under treatment regime $A$ is observed and $T=T^{\prime}_{A}$ and $D=1$.
As ignoring the censoring process biases the likelihood, we assume, as commonly done in survival analysis, that this process is uninformative, formalised as:
Assumption 5 (Non-informative censoring).
Under the previous assumptions, we can maximise the overall likelihood by optimising for the following weighted factual log-likelihood $l_{F}$:
$l_{F}=$ $\sum_{i,d_{i}=1}w_{i}\log\left(\frac{\partial S(t\mid x_{i},a_{i})}{\partial t}\bigg{|}_{t=t_{i}}\right)$
$+\sum_{i,d_{i}=0}w_{i}\log S(t_{i}\mid x_{i},a_{i})$ (2)
where $i$ is the patient index, $(x_{i},t_{i},a_{i},d_{i})$ are realisations of the associated variables $X,T,A$ and $D$, and $w_{i}$ is the inverse propensity weighting correction for patient $i$. Under Assumption 2, the survival corresponds to the weighted survival over the different subgroups:
$S(t\mid x,a):=\mathbb{P}(T^{\prime}\geq t\mid x,a)$
$=\sum_{k=1}^{K}\mathbb{P}(Z=k\mid x,a)\mathbb{P}(T^{\prime}\geq t\mid Z=k,a)$ (3)
In summary, Figure 2 describes all variables and dependences assumed in the studied problem with a directed acyclic graph. Given the observed $X$, $A$, $T$ and $D$, our aim is to estimate the latent structure $Z$ and the associated survival distributions under the two potential treatment regimes $T^{\prime}_{0}$ and $T^{\prime}_{1}$.
Figure 2: Graphical representation between covariates ($X$), treatment ($A$) and outcomes ($T,D$). Realisations of dashed variables are unobserved, while $X$, $A$, $T$ and $D$ are observed.
@@
Figure 3: Causal Survival Clustering (CSC) architecture. Latent parameter $u_{k}$ characterising the subgroup $k$ is inputted in the monotonic network $M$ to estimate the survival under both treatment regimes. $G$ assigns the probability of belonging to each subgroup given the patient’s covariate(s) $x$. To tackle the challenge of treatment assignment bias, the network $W$ estimates the treatment propensity used to weigh the training likelihood. In this context, patient survival is the average of the weighted survival distributions across subgroups under the given treatment.
3.2 Estimating the quantities of interest